OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x – 4| + 1 = 8 Step 2: |x – 4| = 8 – 1 Step 3: |x – 4| = 7 Which of the following is a correct next step to solve the equation?

OpenStudy (anonymous):

x – 4 = –7 x + 4 = –7 –x – 4 = 7 x + 4 = 7

OpenStudy (anonymous):

i think its between A. & C.

OpenStudy (anonymous):

Same

OpenStudy (anonymous):

i think its A.

OpenStudy (anonymous):

no i mean C.

OpenStudy (texaschic101):

the next step has to be, either : (x - 4) = 7 or -(x - 4) = 7

OpenStudy (anonymous):

so it is C.

OpenStudy (texaschic101):

I suppose...I would have thought they would have included the parenthesis

OpenStudy (texaschic101):

yes....I think I would go with C

OpenStudy (anonymous):

I went with c